Nearest point theorem: letF be a non-void closed subset of Rp and let x be a point outside of F. Then there exists at least one point y belonging to F such that ||z - x|| greater than or equal to ||y -x|| for all z in F.

Given the theorem, answer the question:

Does the nearest point theorem in R imply that there is a strictly positive real number nearest zero?
Answer #1

Let d = inf{||2 – 3|| : IEF ​​​​​​.Since F is nonempty, by the definition of infimum for all neN, there exists In EF, such that + P> ||42 - 3|| SP . Hence we can find a sequence {n} CF , such that + P> ||42 - 3|| SP . Thus we have for all neN,  1|2n|| = || In - I+x|| < ||2 - In|| + ||0|| < d + ||0|| + = <d+ ||2|| +1. Hence we have a bounded sequence {n} CF . Then by Bolzano–Weierstrass theorem every bounded infinite set has a limit point, say y, such that d= 11 - y. Note that y is also a limit point of F, by the construction of \{x_n\}. Since F is closed we have Y EF.

Now we have d=||x-y||=\inf_{z\in F}\{||z-x||:x\in F\}\le ||z-x|| , for all ZEF.

No the above theorem does not imply there exists strictly positive real number nearest to 0, as if possible there exists such p>0, such that p is nearest to 0, then consider < , which contradicts that p is the nearest to 0.

Feel free to comment if you have nay doubts. Cheers!